# TriSAM: Tri-Plane SAM for zero-shot cortical blood vessel segmentation in VEM images

### v3

@@ Abstract
Figure 1: Comparison of imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can capture sub-micron resolution of blood vessels in the cortex, but are unable to provide the ultrastructure details. (b) Volume electron microscopy (VEM) can achieve a higher resolution to show unbiased details of the vasculature including all types of cells with details.
@@ 2 Related Work
Blood vessel segmentation methods. Most existing VEM image segmentation algorithms were developed for neurons [18] and synapses [19, 20, 21]. The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11]. This marks a significant milestone in connectomics research. The blood vessels were segmented manually in the human cortex [12] and automatically using fully convolutional neural network (FCN) [22] or 3D U-Net [11, 23, 24]. Due to the burden of annotation, efforts have been made to decrease the need for annotations [25]. For images acquired using other related imaging methods, such as light microscopy and MRI, many methods were developed and covered by some reviews [9, 10]. Briefly, the blood vessels could be enhanced based on the tubular nature using filters like diffusion and ridge filtering. Yet, the segmentation of cortical blood vessels in VEM images is seldom explored due to the absence of publicly available benchmarks.
@@ 3.1 Dataset Description
Sample information. We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [11], superior temporal gyrus from an adult macaque [42], and temporal lobe from an adult human [12]. Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details.
@@
Resolution standardization. We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and the nature of capillary diameter. Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content. As shown in Fig. 2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
3.2 Dataset Annotation
[…]
Semi-automatic segmentation error correction. We used the VAST lite software [43] which can accelerate the proofreading process by using provided segmentation results as the drawing or filling template instead of delineating segments manually.
Quality control. We used 3D visualizations and skeleton analysis to detect remaining segmentation errors. Each volume was proofread by two neuroscience experts in multiple rounds until no disagreement.
Figure 3: Dataset statistics: the histogram of the blood vessel radius.
@@ 3.3 Dataset Statistics
We summarize the dataset details in Tab. 1. The BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [44] and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume. The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3 where the dimension is in the “zyx” order.
Figure 4: Problem formulation and challenges. (a) The 3D segmentation task can be formulated as the multi-seed object tracking problem. There are two main challenges for such an approach: (b) generating high-quality initial seeds (blue dots) and minimizing the false positive (red crosses) and false negative errors (yellow circles); (c) adaptively selecting tracking directions as the quality of SAM segmentation results can vary significantly.
Figure 5: TriSAM framework. (a) Tri-plane selection is first proposed to select the best plane for tracking. (b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking direction. (c) Recursive seed sampling exploits potential turning points for long-term tracking.
@@ 4.1 Overview
We formulate the 3D blood vessel segmentation task as a multi-seed video object segmentation task: generating one 3D segment per seed through tracking 2D SAM results along one fixed direction and the final output is the connected component of the combined segmentation results (Fig. 4a). However, unlike videos, the 3D blood vessel segment can be tracked from different axes, whereas 2D SAM segmentation results can be of drastically different qualities. As shown in Fig. 4b, the seed generation can have many errors leading to incomplete segmentation results. To achieve long-term tracking, the framework needs a recursive seed sampling module to propose potential turning points as new seeds based on segmentation results. As shown in Fig. 4c, the ground truth segment on some 2D planes can have such an unnatural shape that SAM fails to segment it, not to mention its drastic shape change along the axis. Ideally, one axis is along the blood flow direction where the 2D tracking plane is an ellipse shape, e.g., zx plane in the example. Thus, it is essential to have a tri-plane selection module in the framework.
The proposed TriSAM framework has the following three modules: tri-plane selection, SAM-based tracking, and recursive seed sampling (Fig. 5).
4.2 Tri-Plane Selection
@@
Rather than videos, 3D VEM images can be interpreted in different planes. In practice, tracking along some of the planes is hard since the shape and scale of the blood vessels change dramatically. This phenomenon can be further confirmed by the experiment shown in Table 3 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis. Given the seed position as a prompt, we run SAM on each of the tri-planes of the 3D image volume centered at the seed. Note that the plane that is most tangent to the blood flow direction has the smallest segment size, capturing the blood vessel’s cross-section (Fig. 5a, xy-plane). In addition, SAM outputs the probability for the segment result, reflecting the confidence of the naturalness of its shape. Thus, to combine the above cues for tri-plane selection, we pick the 2D plane with the smallest segment size and a probability higher than the threshold $\tau$ (Select in Alg. 1). TriSAM terminates when no plane is selected.
4.3 SAM-based Tracking
Given the segment on the selected 2D plane from above, we aim to track it with SAM along the orthogonal axis (SAMBasedTrack in Alg. 1). If we use the center of the current segment as the point prompt for the next slice, SAM may output segments with inconsistent sizes along the axis due to the lack of size information in the point prompt. If we use the bounding box of the current segment as the box prompt, SAM may not track in the long term as the output enclosed in the bounding box is getting smaller and smaller. Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results. TriSAM terminates when SAM returns no prediction for the next image slice (probability $<\tau$ ).
4.4 Recursive Seed Sampling
The above modules work well when the blood vessel travels in one direction. In reality, blood vessels not only turn in different directions but also bifurcate, leading to drastic shape and scale changes as shown in Fig. 11. Thus, we need a module to propose potential turning points to recursively run TriSAM, linking short-term segment tracks into long-term segment tracks. SAM-based tracking is simple and fast for short-term tracking. But it may fail when shape and scale change dramatically as shown in Fig. 11. Therefore, it relies on the large number of seeds which is hard to generate as shown in Fig. 4(b).
To enable long-term tracking, we propose a recursive seed sampling method to exploit the 3D blood vessel structure, which keeps track of potential turning points in blood vessels (RecursiveSample in Alg. 1). Given the 3D segment from the short-term track after TriSAM terminates, we first use the Farthest Point Sampling (FPS) to find potential endpoints. Then, we add these points to the global seed tree whose root note is the input seed (LeafSplit in Alg. 1). To avoid repeating the track along the same direction as the previous short-term track, these newly sampled seeds store the previous tracking direction. For example, if the previous track is along the z-axis, TriSAM on the new seed will not compute results along the z-axis to encourage exploration of new tracking directions (Valid in Alg. 1). During each iteration, we take an unvisited node in the seed tree, mark it as visited, and run TriSAM along a different tracking direction. With such a seed tree data structure, we run TriSAM recursively until all potential points are visited. For each endpoint, we re-run the above two TriSAM modules to track segments along directions that are different from the current tracking axis.
Mathematically, we denote the 3D coordinate of a turning point as ($x_{t}$,$y_{t}$,$z_{t}$). At this point, we perform the SAM operation on three planes separately. We use $M_{z}(x,y,z_{t})$, $M_{y}(x,y_{t},z)$, and $M_{x}(x_{t},y,z)$ to represent the segmented mask region along xy plane, xz plane, and yz plane, respectively. Assuming that the chosen best plane is xz plane which means we will track along the y-axis, we will obtain new turning points based on the segmented masks on the non-best planes $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$. Since the boundary region of the mask corresponds to where blood vessels disappear, it indicates locations where there is a change in the direction of blood vessel flow. So, the boundaries in the y-direction of both $M_{z}(x,y,z_{t})$ and $M_{x}(x_{t},y,z)$ indicate locations where blood vessels change direction along the y-axis. We further define the maximum and minimum values of the y-coordinate within the $M_{z}(x,y,z_{t})$ region as $y_{max}^{z_{t}}$ and $y_{min}^{z_{t}}$, respectively. Similarly, we can define $y_{max}^{x_{t}}$ and $y_{min}^{x_{t}}$. Then the new turning points can be defined as $\{(x,y,z)|(x,y,z)\in M_{z}(x,y_{max}^{z_{t}},z_{t})\cup M_{z}(x,y_{min}^{z_{t}},z_{t})\cup M_{x}(x_{t},y_{max}^{x_{t}},z)\cup M_{x}(x_{t},y_{min}^{x_{t}},z)\}$. The turning point will be dropped if it hits the volume boundary or the probability of its segment is less than the threshold $\tau$. After updating the turning point tree, we then sample a new leaf node for the next iteration until all leaf nodes are sampled.
In practice, we maintain a tree of seed points where the root is the initial seed. In each iteration, we sample a left node in the tree, mark it as visited, and run TriSAM. Empirically, we find the non-best planes potentially identify turning points or endpoints accurately. Therefore, we propose to generate the turning points according to the segmentation results at non-best planes. Potential turning points are identified as points in these segments with the smallest and largest z-values. The algorithm of TriSAM is shown in Algorithm 1.
Algorithm 1 TriSAM algorithm
[…]
Initialize the seed tree $SeedTree=\{s\}$
for ${\textnormal{t}}_{i}\quad in\quad\operatorname{RemainNode}(SeedTree)$ do
@@ 4.4 Recursive Seed Sampling
     if $\operatorname{Valid}(Seeds)$ then
         $SeedTree=\operatorname{LeafSplit}(SeedTree,{\textnormal{t}}_{i},Seeds)$triangleright Grow the seed tree
@@ 5.1 Experimental Settings
Implementation Details. Since the volume is too large, non-overlapping subvolumes are used for segmentation and the results are fused to form the final prediction. Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background. Therefore, pixels darker than a threshold are considered as background while others are seed candidates. To improve efficiency, we only keep the center of each connected component as the final seeds. For BvEM-Macaque, we directly predict the whole volume. For others, $k\times 1024\times 1024$ subvolumes are used since it is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse. For preprocessing, we utilize deflickering method and temporal smoothing on the z-axis to remove the artifacts and missing slices. For postprocessing, holes are filled and small connected components are removed after the segmentation. Global color thresholding is used to generate the initial seeds and $\eta$ is set to 0.98 as shown Fig. 8 (left). Given the significantly greater complexity of human blood vessels, we employ the manually generated seeds for the challenging BvEM-Human volume. For SAM-based tracking, the probability threshold $\tau$ is set to 0.8. according to the experiment shown in Fig. 8 (right). Unless explicitly stated otherwise, we use MobileSAM [35] instead of the standard SAM [15] in all conducted experiments to improve the inference speed. All experiments are conducted on an NVIDIA-A100 GPU.
@@ 5.2 Benchmark Results
Methods in Comparison. We compare the proposed TriSAM with both zero-shot baselines and supervised methods. The compared zero-shot methods are global color thresholding and SAM+IoU tracking. For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10x512x512). Then we label all voxels that are 3 standard derivations above mean as positive. Finally, connected components with less than 1000 voxels are filtered out. For SAM+IoU Tracking, we segment all objects in each z-slice of the dataset using automatic mask generation. Then we track each blood vessel using the first labeled slice as seeds. Our simple tracking algorithm finds the mask in the next slice with maximum IoU with the current slice. If the max IoU is above a threshold, we assign this mask to the current object and continue tracking. We also tried SAM+IoU tracking with microSAM [48] weights that have been finetuned on EM images. This model however does not work well on our dataset. We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution dataset. We further compare TriSAM with the supervised method 3D U-Net [11]. We use the implementation from [49]. For supervised methods, we cropped subvolumes at the center of each volume composing approximately 10% in size. These annotated subvolumes were divided into a 1-1 train-val split.
@@
5.3 Ablation Studies
Effect of the Tri-plane Selection. We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque volume are shown in Tab. 3. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.
@@ 5.3 Ablation Studies
Figure 7: Segmentation results with different plane selection strategies.
Figure 8: Ablation study results on hyperparameters $\eta$ and $\tau$.

### v4

@@ Abstract
Figure 1: Imaging modalities for blood vessel analysis. (a) Both microtomography ($\mu$CT) [1] and light microscopy (LM) [2] can only capture blood vessels in the cortex at the sub-micron resolution without ultrastructure details. (b) Volume electron microscopy (VEM) can show unbiased details of the vasculature including all types of cells at a higher resolution.
@@ 2 Related Work
Blood vessel segmentation methods. Most existing VEM image segmentation algorithms were developed for neurons [18] and synapses [19, 20, 21]. The traditional VEM dataset size is too small to study blood vessel architecture. Recently, with rapid technology improvement, VEM sample size reached the cubic millimeter scale covering all the layers of cerebral cortex [12, 11], providing rich details of vasculature structure at scale. However, due to the lack of efficient and effective automatic methods, the blood vessels were segmented manually in the human cortex dataset [12]. On another front, automatic blood vessel segmentation methods have been developed for other image modalities, such as light microscopy and MRI, at a lower resolution. We refer to  [9, 10] for detailed reviews on filter-based methods [22, 23] and deep learning methods [24, 25]. Due to the burden of annotation, efforts have been made to decrease the need for annotations [26]. In this paper, we propose the first VEM cortical blood vessel benchmark to foster novel methods.
@@ 3.1 Dataset Description
Tissue samples. We built the BvEM dataset upon the largest publicly available VEM image volumes for three mammal species: visual cortex from an adult mouse [11], superior temporal gyrus from an adult macaque [43], and temporal lobe from an adult human [12]. Each volume was acquired with different protocols at different facilities and we refer to respective papers for more details. As shown in Fig. 2, the imaging quality and the appearance of blood vessels vary drastically across these three volumes, showcasing the cutting-edge large-scale VEM imaging pipelines.
Image volumes. We processed the original image volumes into a standardized form that is suitable for benchmarking. We first downsampled all three VEM image volumes to a near-isotropic resolution ($\sim$200-300 nm) along each dimension, which is a good balance between rich image details for biological analysis and the nature of capillary diameter. Then, we trimmed the image boundary of the BvEM-Mouse and BvEM-Human volumes, where the blood vessels are hard to annotate due to the missing image content.
3.2 Dataset Annotation
[…]
Automatic error detection. For the false split errors, which are the majority source of error, we computed the skeleton of the segmentation and detected all the skeleton endpoints that do not touch the volume boundary as candidates. For the false merge errors, we computed the intersection-over-union (IoU) of 2D segments on neighboring slices and detected areas of small IoU value as candidates.
Manual proofreading. We used the VAST lite software [44] to accelerate the manual proofreading process by using the provided dense segmentation results as templates. Instead of manually delineating segments manually, proofreaders can coarsely draw or fill the segment mask that is snapped to the detailed boundary in the template. Assisted with the 3D blood vessel visualizations, two neuroscience experts proofread each volume in multiple rounds until no disagreement.
@@ 3.3 Dataset Statistics
As shown in Tab. 1, the BvEM-Macaque volume has around 0.5G voxels, and the BvEM-Mouse and BvEM-Human volumes are around 80 and 121 times bigger, respectively. From the blood vessel instance segmentation annotation, we automatically extracted skeleton centerlines [45] and computed the length for each blood vessel instance. Due to the hyper-connectivity nature of cortical blood vessels, the length of the largest instance is around 99%, 95%, and 85% for each volume. The histogram of the blood vessel radius in the proposed dataset is shown in Figure 3. The mouse and the human dataset have similar radius distributions peaked around capillaries, while the macaque volume has no capillaries.
Figure 3: Dataset statistics. Empirical distribution of blood vessel radius.
Figure 4: TriSAM framework. (a) Tri-plane selection is first proposed to select the best plane for tracking. (b) SAM-based tracking leverages SAM to perform short-term tracking given a seed location and a tracking axis. (c) Recursive seed sampling exploits potential turning points for long-term tracking.
@@ 4.1 Overview
We aim to leverage the 2D generalist Segment Anything Model (SAM) to build a zero-shot 3D VEM blood vessel segmentation method, which can be widely applied to various image volumes. Given an input seed position, existing SAM-based tracking methods work well when the blood vessel travels only in one axis. However, in reality, blood vessels not only turn in alternating directions but also bifurcate, leading to drastic shape and scale changes.
Instead, we represent the whole connected blood vessel segment that contains the input seed as a graph, $G=(V,E)$, where each node $v\in V$ is a tubular branch segment that does not bifurcate. Thus, the segmentation problem can be solved with the graph traversal algorithm. After visiting one node (i.e., running SAM-based tracking methods given the seed), we sample new seed positions to find neighboring nodes to segment recursively.
Thus, the proposed TriSAM framework has the following three modules: (1) Tri-plane selection module to pick the best 2D plane to segment the object (Fig. 4a); (2) SAM-based tracking module to predict 2D masks along the selected axis (Fig. 4b); (3) Turning point sampling module to propose potential turning points as new seeds based on existing segmentation results (Fig. 4c) to recursively grow all branches of the blood vessel. Note that the second module can be any existing SAM-based tracking method.
4.2 TriSAM Framework
We first explain the three modules in the TriSAM framework (Fig. 4) and then provide the algorithmic description (Alg. 1). In practice, we sample many initial seeds, run the TriSAM framework for each seed, and fuse the segmentation results.
Tri-plane selection module. Unlike videos, 3D VEM image volumes $I$ can be segmented on 2D planes along different axes, $p\in\{x,y,z\}$, based on the object morphology. However, tracking along some axis $p$ can be harder than others due to the irregular shape of the 2D cross-section of the 3D blood vessels on the corresponding 2D plane $I_{p}$. We later verify such a phenomenon empirically as shown in Table 3 in which tracking along the $y$ axis is more effective than tracking along the $x$ axis for a given seed position.
@@ 4.2 TriSAM Framework
This module aims to select the best axis $p^{*}$ for tracking, given the SAM segmentation results, $\mathrm{SAM}(I_{p},s)$, on each tri-plane $I_{p}$ centered at the seed position $s$ as the approximation of the blood vessel cross-sections. In theory, the ideal axis for tracking is along the blood flow direction, where the corresponding 2D image plane is tangent to the blood flow direction with the smallest cross-section segment area (Fig. 4a, xy-plane). In addition, SAM outputs the probability, $P_{\mathrm{SAM}}$, for the segment result, reflecting the confidence in the naturalness of its shape. Thus, to combine the two cues above, we pick the axis with the smallest segment size on the 2D plane and a probability of at least the threshold $\tau$ (plane-select in Alg. 1).
$p^{*}=\arg\min_{p}\mathrm{Area}(\mathrm{SAM}(I_{p},s)),~{}\mathrm{s.t.}~{}P_{\mathrm{SAM}}(I_{p},s)>\tau.$ (1)
To avoid an infinite recursion, traditional graph traversal algorithms check if a node is marked visited. In our case, a tubular branch segment may grow into different blood vessel segments along different tracking axes, e.g., bifurcation regions. Thus, we mark the combination of the seed and the tri-plane tracking axis $(s,p^{*})$ visited if seed $s$ falls in the segment that is predicted through tracking along $p^{*}$ axis (Visited in Alg. 1). Note that $p^{*}=\phi$ if no plane is selected and Visited($(s,p^{*})$) is True. To implement the “Visited” function, we store both the segmentation result and the $p^{*}$, which is omitted in Alg. 1 for simplicity.
SAM-based tracking module. Given the seed position $s$, the selected tracking axis $p^{*}$ and SAM’s initial 2D segment mask, we need to produce a 3D segment by tracking the 2D mask with SAM in both directions (e.g., $x+$ and $x-$). Note that naively propagating the mask center or bounding box as the SAM prompt for the next image slice leads to poor segmentation results, as SAM may output segments with inconsistent sizes or shrinking sizes along the axis respectively. Many existing sophisticated SAM-based tracking methods can be directly used.
This module aims to provide a simple yet effective approach by generating better SAM prompts for the next image slice. Empirically, we find prompting SAM with both the enlarged bounding box and the center of the segment generates better segment tracking results (SAM-Track in Alg. 1).
$prompt=\{(x,y),(x,y,\gamma w,\gamma h)\},$ (2)
where $(x,y,\gamma w,\gamma h)$ is the bounding box of the segmentation and $\gamma$ is the scaling factor. To ensure the quality of the predicted SAM mask, the proposed tracking module terminates when the SAM probability for the predicted mask is lower than the threshold $\tau$.
Turning point sampling module. As discussed above, SAM-based tracking module can not handle blood vessel segments that change directions or bifurcate, where the 2D segment along the original direction changes drastically. Thus, when the SAM-based tracking module terminates in the middle of the volume, we need to find turning points as new seeds to track segments along other directions. Naively, we can densely sample points from the segmentation result and run the plane-select module to find points that prefer other directions. However, such an approach can be inefficient as most points may prefer the original tracking direction, and ineffective due to false turning points caused by SAM errors.
Instead, we design this module to sample turning points around the last point prompt position from the SAM-based tracking module. We first predict 2D SAM segmentation $seg_{p}$ along the other two directions ($p\in\{x,y,z\}\setminus\{p^{*}\}$) at the last point prompt position, which has a high probability of capturing the tangent cross-section of the swerving blood vessel. Then, we sample $K$ points from each of the 2D segmentation $seg_{p}$ with the Farthest Point Sampling (FPS) method as turning points added to the seed list.
$seeds=\bigcup_{p\in\{x,y,z\}\setminus\{p^{*}\}}\mathrm{FPS}(seg_{p},K).$ (3)
Algorithm 1 TriSAM Framework
3D image volume $I$ and initial seed $s_{0}$, threshold $\tau$
@@
     $p^{*}=\textsc{plane-select}(I,s)$
     if not $\textsc{Visited}(s,p^{*})$ then
         $seg=\textsc{SAM-Track}(I,s,p^{*})$
         if $seg\neq\phi$ then
              $seeds=\textsc{TurningPoint-Sample}(seg,p^{*})$
[…]
@@
[…]
4.3 Implementation Details
Image Pre-processing. The large-scale datasets have many image artifacts and missing slices, which degrade the SAM-based tracking results. We utilize the temporal smoothing method along the z-axis to deflicker the images.
Initial seed generation. Initial seeds can be effectively generated with global color thresholding since the pixels of blood vessels are brighter than the background. To improve efficiency, we only keep the center of each connected component as the final seeds.
SAM details. Unless stated otherwise, we use MobileSAM [36] instead of the standard SAM [15] in all conducted experiments to improve the inference speed. To clean up SAM results, holes are filled and small connected components are removed through binary morphological operations.
Hyperparamter selection. As a zero-shot approach, TriSAM sets the hyperparameters to reasonable values instead of using ground truth data. In later ablation studies, we verify the robustness of hyperparameters within the reasonable range. For the global color thresholding for seed generation, $\eta$ is set to 98 percentile. For SAM-based tracking, the probability threshold $\tau$ is set to 0.8.
Inference. For BvEM-Macaque, we directly predict the whole volume. For others, $k\times 1024\times 1024$ subvolumes are used since 1024$\times$1024 is the default resolution for SAM where $k=661$ for BvEM-Human and $k=818$ for BvEM-Mouse. The subvolume results are later fused to form the final prediction. All experiments are conducted on an NVIDIA-A100 GPU.
@@ 5.2 Benchmark Results
Methods in Comparison. We compare the proposed TriSAM with both zero-shot baselines and supervised methods. The compared zero-shot methods are global color thresholding and SAM+IoU tracking. For color thresholding, we first perform (3D) Gaussian blurring with $\sigma=1$ on 3D chunks (10$\times$512$\times$512). Then we label all voxels that are 3 standard derivations above mean as positive. Finally, connected components with less than 1000 voxels are filtered out. For SAM+IoU Tracking, we segment all objects in each z-slice of the dataset using automatic mask generation. Then we track each blood vessel using the first labeled slice as seeds. Our simple tracking algorithm finds the mask in the next slice with maximum IoU with the current slice. If the max IoU is above a threshold, we assign this mask to the current object and continue tracking. We also tried SAM+IoU tracking with microSAM [49] weights that have been finetuned on EM images. This model however does not work well on our dataset. We expect this is because microSAM has been finetuned on high-resolution EM images and does not generalize to our low-resolution dataset. We further compare TriSAM with the supervised method 3D U-Net [11]. We use the implementation from [50]. For supervised methods, we cropped subvolumes at the center of each volume composing approximately 10% in size. These annotated subvolumes were divided into a 1-1 train-val split.
@@
Figure 6: Results with oracle 2D segmentation.
Figure 7: Results for different plane selection strategies without recursive seed sampling.
@@
5.3 Ablation Studies
Oracle analysis. We conduct an oracle analysis to showcase the advantage of the proposed TriSAM framework for its capability to deal with complicated vascular geometry. For the SAM-based tracking module in TriSAM, we plug in an oracle 2D segmentation method which returns the connected component of the ground truth mask containing the prompt. As shown in Fig. 7b, the TriSAM framework can perfectly segment the whole vasculature with the oracle 2D segmentation. In comparison (Fig. 7a), existing SAM-based video object tracking methods can only track along one axis, missing bifurcated branches.
Tri-plane selection module. We first compare our method with the single-plane methods to evaluate the effectiveness of the tri-plane. For the single-plane method, we choose one plane as the main plane and only track along the chosen plane. The results on the BvEM-Macaque volume are shown in Tab. 3. We have observed significant differences in performance among the three single-plane methods, with accuracy ranging from 11.53% to 50.75%, depending on the chosen tracking plane, which indicates the importance of the chosen tracking plane. This variability can be attributed to the tubular nature of blood vessel extensions within biological organisms, resulting in the generation of intricate mask shapes in certain planes, while simpler mask shapes are produced in others. Then we fuse the results of three Single-Plane methods and attain a higher accuracy of 57.41%, which demonstrates that the segmentation results from different planes exhibit a high degree of complementarity with each other. Instead, We see that tri-plane selection exploits the blood vessel 3D structures by tracking along a suitable plane and attains the highest accuracy of 62.14%.